# Rejection-Sampled Linear Codes for Lossy Compression and Channel Simulation

### v1

@@ Abstract
We show that a linear code combined with rejection sampling can give a capacity-achieving scheme for simulating channels with additive noises with exchangeable distributions. Hence, it can be used in lossy source coding to achieve the rate-distortion function. Interestingly, unlike conventional linear covering codes for lossy compression which concerns the trade-off between the rate and the covering radius, our construction only requires the linear code to have a large distance (not a large covering radius), and is not sensitive to the rate of the linear code. Experiments reveal that our construction can outperform conventional covering codes for lossy source coding with Hamming distortion for a certain range of distortion levels, and performs well even when the blocklength is small (e.g., 24).
@@ I Introduction
In this work, we propose combining greedy rejection sampling [4] and a linear code (which we call the inner code), yielding an efficient channel simulation scheme for small blocklengths. Our scheme can be applied to any additive noise channel with finite field vector input , i.e., $\mathbf{X}\in\mathbb{F}_{q}^{n}$, $\mathbf{Y}=\mathbf{X}+\mathbf{Z}$, where $\mathbf{Z}$ has an exchangeable distribution. We show that, as long as the inner code has a large enough distance, our scheme uses an amount of communication close to the capacity $C=\max_{p_{\mathbf{X}}}I(\mathbf{X};\mathbf{Y})$, and hence is capacity-achieving.
@@ II-A Channel Simulation
It has been observed by Steiner [21] that rejection sampling [22] can be utilized to construct a one-shot channel simulation scheme. To maximize the acceptance probability at each iteration, Harsha et al. [4] proposed the greedy rejection sampling for discrete distributions. This approach was extended to general distributions by Flamich and Theis [9]. It was shown that the expected length of iteration number, i.e., the expected communication, is lower-bounded by $D_{\mathrm{KL}}(P\|Q)$ and upper bounded by $D_{\mathrm{KL}}(P\|Q)+2\log(D_{\mathrm{KL}}(P\|Q)+1)+O(1)$, where $P$ and $Q$ are the target and proposal distributions, respectively. By introducing a new quantity, the channel simulation divergence $D_{\mathrm{CS}}(\cdot\|\cdot)$, Goc and Flamich [23] further tightened the both bounds to $D_{\mathrm{CS}}(P\|Q)$ and $D_{\mathrm{CS}}(P\|Q)+\log(e+1)$, which implies the lower bound is achievable tightly. They also revealed that that for any causal rejection sampling scheme, the expected time complexity grows at least exponentially in the Rényi $\infty$-divergence $D_{\infty}(P\|Q)$ between $P$ and $Q$. To improve the running time, Flamich et al. [24, 10, 25] proposed an approach to partition the sample space so as to allow more efficient searching for the desired sample. However, this approach requires additional assumptions on $P$ and $Q$ to ensure efficiency. Note that the sampling schemes are unstructured. For practical implementation, it may be preferable to use structured approaches. Based on the capacity-achieving polar codes [26], Sriramu et al. [12] developed a simulation scheme for $n$ copies of symmetric binary-output channels, which shows asymptotic rate-optimality with time complexity $O(n\log n)$. In addition, they proposed using general linear codes with a correction step to simulate binary symmetric channels. Although both schemes approach the capacity for large $n$, their performance in short blocklength regimes is not guaranteed. Refer to [27] for an overview of various channel simulation schemes.
@@ II-C Oblivious Relaying
Another application is the information bottleneck channel, also known as the oblivious relay channel [18, 19, 20], which concerns the setting where the encoder encodes the message $M$ into an input sequence $\mathbf{X}=(X_{1},\ldots,X_{n})$, which is sent through a memoryless channel $p_{Y|X}$. A relay observes the output $\mathbf{Y}$ and compresses it into a description $W$ at a rate $R_{\mathrm{relay}}>0$. The decoder observes $W$ and attempts to decode $M$. The relay is required to be oblivious, that is, it cannot access the codebook used by the encoder and the decoder. The capacity is given by the information bottleneck [40] $\max_{p_{U|Y}:I(Y;U)\leq R_{\mathrm{relay}}}I(X;U)$ where $X\to Y\to U$ forms a Markov chain [18]. For example, consider $p_{Y|X}$ to be a binary symmetric channel $\mathrm{BSC}(\alpha)$. If the relay simulates the channel $p_{U|Y}=\mathrm{BSC}(\beta)$ so as to allow the decoder to recover an auxiliary output $\mathbf{U}$, the overall channel from $\mathbf{X}$ to $\mathbf{U}$ would be $\mathrm{BSC}(\alpha(1-\beta)+(1-\alpha)\beta)$. Hence, the encoder and decode can utilize standard coding schemes for the BSC on the auxiliary channel from $\mathbf{X}$ to $\mathbf{U}$ in order to achieve the capacity, and no specialized algorithm is needed to accommodate the oblivious relay. In comparison, if the relay utilized lossy source coding to compress $\mathbf{Y}$ (guaranteeing only a small Hamming distance between $\mathbf{Y}$ and $\mathbf{U}$) instead of channel simulation, then it would be difficult to compute the likelihood of $\mathbf{X}$ given $\mathbf{U}$. See Figure 1. Refer to [41] for the connection between oblivious relaying and channel simulation.
@@
Logarithm is base $2$, and entropy is in bits. Write :=$\mathbb{N}:=\{1,2,\ldots\}$, $[a:b]:=\{a,\ldots,b\}$, $[n]:=\{1,\ldots,n\}$. For an event $E$, write $\mathbf{1}\{E\}$ for the indicator function which is $1$ if $E$ occurs, 00 otherwise. Write $\mathbb{F}_{q}$ for the finite field of order $q$. All vectors are assumed to be row vectors. For a vector $\mathbf{x}\in\mathbb{F}_{q}^{n}$, its Hamming weight is $\mathrm{wt}(\mathbf{x}):=|\{i\in[n]:\,x_{i}\neq 0\}|$.
Let $\mathcal{P}_{n}(\mathcal{X})$ be the set of probability mass functions over $\mathcal{X}$ where each entry is a multiple of $1/n$. For $\mathbf{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, its type (or empirical distribution) is denoted as
$\mathrm{type}(\mathbf{x}):=(n^{-1}|\{i:\,x_{i}=y\}|)_{y\in\mathcal{X}}\in\mathcal{P}_{n}(\mathcal{X}).$
@@ III The Channel Simulation Setting
In the setting of (one-shot) channel simulation with common randomness [4], there is a common randomness $B\sim p_{B}$ available to both the encoder and the decoder. The encoder observes an input $\mathbf{X}$, and sends a description $M=f(B,\mathbf{X})\in\mathcal{C}_{M}$ to the decoder, where $\mathcal{C}_{M}\subseteq\{0,1\}^{*}$ is a prefix-free codebook. The decoder then outputs $\hat{\mathbf{Y}}=g(B,M)$. The objective is to find a scheme $(p_{B},f,g)$ that ensures that $\hat{\mathbf{Y}}$ follows a prescribed conditional distribution $p_{\mathbf{Y}|\mathbf{X}}$ given $\mathbf{X}$ exactly or approximately, while minimizing the amount of communication $\mathbb{E}[|M|]$ (the expected length of $M$). In other words, we are simulating the noisy channel $p_{\mathbf{Y}|\mathbf{X}}$ using the noiseless communication of $M$. See Figure 2.
@@
  • Constant weight channel, where $\mathbf{Z}\sim\mathrm{Unif}(\{\mathbf{z}:\,\mathrm{wt}(\mathbf{z})=w\})$ for a fixed $w\in[0:n]$. Exactly simulating this channel also ensures the almost-sure distortion constraint.
@@ IV Stochastic Syndrome Coder
  • Syndrome decoder for channel codes. The encoder encodes the message $\mathbf{M}\in\mathbb{F}_{q}^{k}$ into the codeword $\mathbf{X}=f(\mathbf{M})\in\mathbb{F}_{q}^{n}$ satisfying $\mathbf{X}\mathbf{H}^{\top}=\mathbf{0}$. The decoder observes $\mathbf{Y}\in\mathbb{F}_{q}^{n}$, generates $\hat{\mathbf{Z}}\in\mathbb{F}_{q}^{n}$ following the conditional distribution $\kappa(\hat{\mathbf{z}}|\mathbf{Y}\mathbf{H}^{\top})$, computes $\hat{\mathbf{X}}:=\mathbf{Y}-\hat{\mathbf{Z}}$, and finds the unique $\hat{\mathbf{M}}\in\mathbb{F}_{q}^{k}$ with $\hat{\mathbf{X}}=f(\hat{\mathbf{M}})$.
@@
In both settings, the conditional distribution $\kappa$ is often taken to be the deterministic mapping which maps the syndrome $\mathbf{s}\in\mathbb{F}_{q}^{n-k}$ to the coset leader $\mathrm{argmin}_{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}}\mathrm{wt}(\mathbf{z})$.
In this paper, we propose to use the stochastic syndrome coder for channel simulation. Compared to lossy source coding, we include a common randomness $(\boldsymbol{\Pi},\mathbf{b})$ to perform a random permutation and shift, so that the noise $\hat{\mathbf{z}}$ is uniform within any type class (a similar randomization is also present in [12]).
@@
We now verify that the stochastic syndrome encoder simulates an additive exchangeable noise channel.
@@
The stochastic syndrome encoder simulates an additive exchangeable noise channel, with noise distribution
@@
Due to the random permutation, $\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}})$ only depends on $\mathrm{type}(\hat{\mathbf{z}})$. ∎
@@ V Rejection-Sampled Syndrome Encoder
A limitation of the stochastic syndrome encoder is that it is a fixed-length scheme, i.e., the length of the description $\mathbf{M}$ is always $k$. A fixed-length scheme can be suitable for approximate channel simulation where the actual noise simulated $\hat{\mathbf{Z}}$ only approximately follows the distribution of the desired noise $\mathbf{Z}$ (e.g., i.i.d. Bernoulli for binary symmetric channel), but is generally not suitable for exact channel simulation where $\hat{\mathbf{Z}}$ must have the same distribution as $\mathbf{Z}$ [27]. In this section, we will describe a class of variable-length schemes which combine the stochastic syndrome encoder with rejection sampling, suitable for exact simulation.
@@
We call the sequence of codes $(\mathbf{H}_{i})_{i}$ the inner code sequence. Definition 3 describes a general class of schemes instead of one specific scheme. The construction of a specific capacity-achieving RSSE will be discussed in Section VI. Note that each $(\mathbf{H}_{i},\kappa_{i})$ in an RSSE is almost a stochastic syndrome encoder, except that $\kappa_{i}$ is allowed to output a “rejection symbol” $\mathrm{e}$ which signals that we should reject the current iteration and proceed to $(\mathbf{H}_{i+1},\kappa_{i+1})$. To perform encoding in the RSSE, the encoder performs the same steps as in the stochastic syndrome encoder in Section IV on $(\mathbf{H}_{i},\kappa_{i})$ for $i=1,2,\ldots$ (each with an independent common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$) until the first time where $\hat{\mathbf{V}}_{i}\neq\mathrm{e}$ (let it be iteration $L\in\mathbb{N}$), and selects $\hat{\mathbf{Z}}=\hat{\mathbf{V}}_{L}\boldsymbol{\Pi}_{L}^{-1}$, $\hat{\mathbf{Y}}=\mathbf{X}+\hat{\mathbf{Z}}$. The encoder then encodes $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$, and sends $(L,\mathbf{M})$. Practically, the encoder encodes $L$ using the Huffman code [43] or Elias codes [44], and then append $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$ to form the transmitted bit sequence, which is prefix-free since the decoder can recover $L$ and then know the length of $\mathbf{M}$ which is $k_{L}$. The decoder can then decode $\mathbf{M}$ into $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$. If the Huffman code is used, the expected communication cost is at most $H(L)+\mathbb{E}[k_{L}\log q]+1$.
The precise algorithm for RSSE is given in Algorithm 1, where we assume there are two synchronized random number generators (RNG) available to the encoder and the decoder which can be used to generate the common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$. In practice, they will initialize two pseudo-random number generators with the same seed.
@@
7:         $\hat{\mathbf{Z}}\leftarrow\hat{\mathbf{V}}_{i}\boldsymbol{\Pi}_{i}^{-1}$,  $\hat{\mathbf{Y}}\leftarrow\mathbf{X}+\hat{\mathbf{Z}}$
8:         Encode $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$
9:               (e.g., using (1) if $\mathbf{H}_{i}$ is in standard form)
10:         return $(i,\mathbf{M})$
11:     end if
12:end for
@@
    Output: output $\mathbf{X}\in\mathbb{F}_{q}^{n}$
@@
We can check that RSSE simulates an additive exchangeable noise channel.
@@
The rejection-sampled syndrome encoder simulates an additive exchangeable noise channel, with noise distribution

### v2

@@ Abstract
We show that linear codes combined with rejection sampling can yield a capacity-achieving scheme for simulating additive exchangeable noise channels. Specifically, our scheme achieves an amount of communication within $\log e+1$ bits from the excess functional information lower bound. Hence, it can be used in lossy source coding to achieve the rate-distortion function. We discuss practical implementations based on BCH codes and polar codes. For the simulation of binary symmetric channels, the BCH-based construction with a blocklength of $n=63$ attains a rate comparable to the PolarSim with $n=4096$, while significantly reducing the latency. The polar-based construction asymptotically achieves the channel capacity with polynomial average complexity. Furthermore, using the idea from greedy rejection sampling, we propose an algorithm to construct capacity-achieving schemes based on any linear codes. Experiments reveal that our construction can outperform conventional covering codes for lossy source coding with Hamming distortion for a certain range of distortion levels, and performs well even when the blocklength is small (e.g., $n=24$).
@@ I Introduction
In this work, we propose efficient channel simulation schemes by integrating linear codes (which we call the inner codes) with rejection sampling [13]. Our schemes can be applied to exactly simulate any additive exchangeable noise (AXN) channel with finite field vector input, i.e., $\mathbf{X}\in\mathbb{F}_{q}^{n}$, $\mathbf{Y}=\mathbf{X}+\mathbf{Z}$, where $\mathbf{Z}$ has an exchangeable distribution. Notably, our schemes perform well even at short blocklengths. We show that:
  • The rejection-sampled syndrome encoder (RSSE), which combines syndrome coding of linear codes with rejection sampling, can simulate AXN channels. For a subclass termed restricted type channels (RTCs), it can achieve an amount of communication within $\log e+1$ bits from the channel capacity $C=\max_{p_{\mathbf{X}}}I(\mathbf{X};\mathbf{Y})$.
  • The state-dependent RSSE, which integrates the RSSE with a technique termed RTC-decomposition, can simulate any AXN channel using an amount of communication within $\log e+1$ bits from the excess functional information lower bound [6, Proposition 1]
    $\sum_{\mathbf{y}\in\mathbb{F}_{q}^{n}}\int_{0}^{1}\mathbb{P}(p_{\mathbf{Y}|\mathbf{X}}(\mathbf{y}|\mathbf{X})\geq u)\log\mathbb{P}(p_{\mathbf{Y}|\mathbf{X}}(\mathbf{y}|\mathbf{X})\geq u)\,du,$
    which is tighter than channel capacity for finite blocklengths. Practical implementations based on BCH codes [14, 15] and polar codes [16, 17] are discussed. For the simulation of BSCs, the BCH-based construction with a blocklength of $n=63$ attains a rate comparable to the PolarSim [12] with $n=4096$, resulting in a smaller latency while maintaining computational efficiency. The polar-based construction asymptotically achieves the channel capacity with $O(n^{3/2}\log n)$ average complexity, enabling scalable high-performance simulation.
  • The greedy RSSE (GRSSE) is a construction inspired by greedy rejection sampling [4] for simulating any AXN channel. As long as the inner code has a sufficiently large distance, the constructed RSSE achieves an amount of communication close to the channel capacity $C$, and hence is capacity-achieving.
@@ II-A Channel Simulation
Steiner [26] first observed that rejection sampling [13] can be used to construct channel simulation schemes. To maximize the acceptance probability at each iteration, Harsha et al. [4] proposed greedy rejection sampling (GRS) for discrete distributions. Flamich and Theis [9] and Flamich et al. [27] later extended it to general distributions. Let $L$ denote the number of iterations (samples) required by the GRS with target distribution $P$ and proposal distribution $Q$. It was shown [4] that $H(L)$ (which reflects the expected communication cost) is bounded between $D_{\mathrm{KL}}(P\|Q)$ and $D_{\mathrm{KL}}(P\|Q)+2\log(D_{\mathrm{KL}}(P\|Q)+1)+O(1)$. By introducing the channel simulation divergence $D_{\mathrm{CS}}$, which generalizes the excess functional information lower bound [6, Proposition 1], Goc and Flamich [28] tightened both bounds to $D_{\mathrm{CS}}(P\|Q)\geq D_{\mathrm{KL}}(P\|Q)$ and $D_{\mathrm{CS}}(P\|Q)+\log(e+1)\leq D_{\mathrm{KL}}(P\|Q)+\log(D_{\mathrm{KL}}(P\|Q)+1)+\log(e+1)+1$, showing their lower bound is tight. They also revealed [28] that for any causal rejection sampling scheme, the expected time complexity $\mathbb{E}[L]$ grows at least exponentially in the Rényi $\infty$-divergence $D_{\infty}(P\|Q)$. Poisson functional representation (PFR) [6] provides a noncausal sampling scheme for general distributions. It achieves an expected communication cost upper-bounded by [29]
$I(X;Y)+\log(I(X;Y)+2)+3.$ (1)
To accelerate the simulation process, Flamich et al. [10, 27, 30] proposed partitioning the sample space for more efficient searching. The sampling schemes discussed so far are unstructured. For practical implementation, structured codes are often preferable. Based on the capacity-achieving polar codes [16], Sriramu et al. [12] developed a simulation scheme for $n$ copies of symmetric binary-output channels, which is asymptotically rate-optimal, in the sense that the communication rate approaches the channel capacity, with $O(n\log n)$ complexity. They also proposed using general linear codes with a correction step to simulate BSCs. Although both schemes approach the capacity for large $n$, their performance at short blocklengths is not guaranteed.
Refer to [31] for an overview of various channel simulation schemes.
@@ II-C Oblivious Relaying
Another application is the information bottleneck channel, also known as the oblivious relay channel [23, 24, 25], which concerns the setting where the encoder encodes the message $M$ into an input sequence $\mathbf{X}=(X_{1},\ldots,X_{n})$, which is sent through a memoryless channel $p_{Y|X}$. A relay observes the output $\mathbf{Y}$ and compresses it into a description $W$ at a rate $R_{\mathrm{relay}}>0$. The decoder observes $W$ and attempts to decode $M$. The relay is required to be oblivious, that is, it cannot access the codebook used by the encoder and the decoder. The capacity is given by the information bottleneck [43] $\max_{p_{U|Y}:I(Y;U)\leq R_{\mathrm{relay}}}I(X;U)$ where $X\to Y\to U$ forms a Markov chain [23]. For example, consider $p_{Y|X}$ to be a binary symmetric channel $\mathrm{BSC}(\alpha)$. If the relay simulates the channel $p_{U|Y}=\mathrm{BSC}(\beta)$ so as to allow the decoder to recover an auxiliary output $\mathbf{U}$, the overall channel from $\mathbf{X}$ to $\mathbf{U}$ would be $\mathrm{BSC}(\alpha(1-\beta)+(1-\alpha)\beta)$. Hence, the encoder and decoder can utilize standard coding schemes for the BSC on the auxiliary channel from $\mathbf{X}$ to $\mathbf{U}$ in order to achieve the capacity, and no specialized algorithm is needed to accommodate the oblivious relay. In comparison, if the relay utilized lossy source coding to compress $\mathbf{Y}$ (guaranteeing only a small Hamming distance between $\mathbf{Y}$ and $\mathbf{U}$) instead of channel simulation, then it would be difficult to compute the likelihood of $\mathbf{X}$ given $\mathbf{U}$. See Figure 1. Refer to [44] for the connection between oblivious relaying and channel simulation.
@@
Logarithm is base $2$, and entropy is in bits. Write $\mathbb{N}:=\{1,2,\ldots\}$, $a:b]:=\{a,\ldots,b\}$, $n]:=\{1,\ldots,n\}$. For an event $E$, write $\mathbbm{1}\{E\}$ for the indicator function which is $1$ if $E$ occurs, 0 otherwise. Write $\mathbb{F}_{q}$ for the finite field of order $q$. All vectors are assumed to be row vectors. For a vector $\mathbf{x}\in\mathbb{F}_{q}^{n}$, its Hamming weight is $\mathrm{wt}(\mathbf{x}):=|\{i\in[n]:\,x_{i}\neq 0\}|$.
Let $p_{X}$ denote the probability mass function of a random variable $X$ over $\mathcal{X}$. The support of $p_{X}$ is defined as $\mathrm{supp}(p_{X}):=\{x\in\mathcal{X}:\,p_{X}(x)>0\}$. The uniform distribution over $\mathcal{X}$ is denoted as $\mathrm{Unif}(\mathcal{X})$. Let $\mathcal{P}_{n}(\mathcal{X})$ denote the set of probability mass functions over $\mathcal{X}$ with entries being multiples of $1/n$. For $\mathbf{x}\in\mathcal{X}^{n}$, its type (empirical distribution) is defined as $\mathrm{type}(\mathbf{x}):=(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{x_{i}=x\})_{x\in\mathcal{X}}\in\mathcal{P}_{n}(\mathcal{X})$. The number of sequences in $\mathcal{X}^{n}$ of type $\mathbf{t}=(t_{1},...,t_{|\mathcal{X}|})\in\mathcal{P}_{n}(\mathcal{X})$ is given by the multinominal coefficient
$\binom{n}{n\mathbf{t}}:=\frac{n!}{(nt_{1})!\cdots(nt_{|\mathcal{X}|})!}.$
@@ III The Channel Simulation Setting
In the setting of (one-shot) channel simulation with common randomness [4], there is a common randomness $B\sim p_{B}$ available to both the encoder and the decoder. The encoder observes an input $\mathbf{X}$, and sends a description $M=f(B,\mathbf{X})\in\mathcal{C}_{B}$ to the decoder, where $\mathcal{C}_{B}\subseteq\{0,1\}^{*}$ is a prefix-free codebook. The decoder then outputs $\hat{\mathbf{Y}}=g(B,M)$. The objective is to find a scheme $(p_{B},f,g)$ that ensures that $\hat{\mathbf{Y}}$ follows a prescribed conditional distribution $p_{\mathbf{Y}|\mathbf{X}}$ given $\mathbf{X}$ exactly or approximately, while minimizing the amount of communication $\mathbb{E}[|M|]$ (the expected length of $M$). In other words, we are simulating the noisy channel $p_{\mathbf{Y}|\mathbf{X}}$ using the noiseless communication of $M$. See Figure 2.
@@
  • Constant type channel, where $\mathbf{Z}$ is uniform over a type class $\mathbb{F}_{q}^{n}\{\mathbf{t}\}:=\{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathrm{type}(\mathbf{z})=\mathbf{t}\}$ for a fixed $\mathbf{t}\in\mathcal{P}_{n}(\mathbb{F}_{q})$. If $q=2$, the channel is equivalent to a constant weight channel, where $\mathbf{Z}$ is uniform over $\mathbb{F}_{2}^{n}\{w\}:=\{\mathbf{z}\in\mathbb{F}_{2}^{n}:\,\mathrm{wt}(\mathbf{z})=w\}$ for a fixed $w\in[0:n]$. Exactly simulating this channel also ensures the almost-sure distortion constraint.
  • Restricted type channel (RTC), where $\mathbf{Z}$ is uniform over $\mathbb{F}_{q}^{n}\{\mathcal{T}\}:=\cup_{\mathbf{t}\in\mathcal{T}}\mathbb{F}_{q}^{n}\{\mathbf{t}\}$ for a fixed set of types $\mathcal{T}\subseteq\mathcal{P}_{n}(\mathbb{F}_{q})$. This channel generalizes the Hamming ball and the constant type channels.
@@ IV Stochastic Syndrome Coder
[…]
@@
A common instance in both settings is the minimum distance syndrome coder, where $\kappa$ is taken to be the deterministic mapping which maps the syndrome $\mathbf{s}\in\mathbb{F}_{q}^{n-k}$ to the coset leader ${\arg\min}_{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}}\mathrm{wt}(\mathbf{z})$. Another notable instance is the bounded-distance decoder for channel coding, where $\kappa$ satisfies $\kappa(\mathbf{z}|\mathbf{z}\mathbf{H}^{\top})=1$ for all $\mathbf{z}\in\mathbb{F}_{q}^{n}$ with $\mathrm{wt}(\mathbf{z})\leq\tau$ for a certain error-correction radius $\tau\leq\lfloor(d-1)/2\rfloor$, and $d$ is the minimum distance of the linear code defined by $\mathbf{H}$. Note that the conventional bounded-distance decoder does not specify the output of $\kappa$ when the input syndrome $\mathbf{s}\notin\{\mathbf{z}\mathbf{H}^{\top}:\,\mathbf{z}\in\mathbb{F}_{q}^{n},\mathrm{wt}(\mathbf{z})\leq\tau\}$. To match Definition 1, we assume in such cases $\kappa$ outputs an arbitrary $\mathbf{z}\in\mathbb{F}_{q}^{n}$ satisfying $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$. For example, if $\mathbf{H}=[\tilde{\mathbf{H}}\,|\,\mathbf{I}_{n-k}]$ is in standard form, $\mathbf{z}$ can be $\mathbf{m}\,|\,\mathbf{s}-\mathbf{m}\tilde{\mathbf{H}}^{\top}$] for any $\mathbf{m}\in\mathbb{F}_{q}^{k}$.
In this paper, we propose to use the stochastic syndrome coder for channel simulation. Compared to lossy source coding, we include a common randomness $(\boldsymbol{\Pi},\mathbf{B})$ to perform a random permutation and shift, so that the induced noise $\hat{\mathbf{Z}}$ is uniform within any type class (a similar randomization is also present in [12]).
@@
We now verify that the stochastic syndrome encoder simulates an AXN channel.
@@
The stochastic syndrome encoder simulates an AXN channel, with noise distribution
@@
[…]
@@ V Rejection-Sampled Syndrome Encoder
A limitation of the stochastic syndrome encoder is that it is a fixed-length scheme, i.e., the length of the description $\mathbf{M}$ is always $k$. A fixed-length scheme can be suitable for approximate channel simulation where the induced noise $\hat{\mathbf{Z}}$ only approximately follows the distribution of the target noise $\mathbf{Z}$ (e.g., i.i.d. Bernoulli for binary symmetric channel), but is generally not suitable for exact channel simulation where $\hat{\mathbf{Z}}$ must have the same distribution as $\mathbf{Z}$ [31]. In this section, we will describe a class of variable-length schemes which combine the stochastic syndrome encoder with rejection sampling, suitable for exact simulation.
@@
We call the sequence of codes $(\mathbf{H}_{i})_{i}$ the inner code sequence. Definition 3 describes a general class of schemes instead of one specific scheme. Note that each $(\mathbf{H}_{i},\kappa_{i})$ in an RSSE is almost a stochastic syndrome encoder, except that $\kappa_{i}$ is allowed to output a “rejection symbol” $\mathrm{e}$ which signals that we should reject the current iteration and proceed to $(\mathbf{H}_{i+1},\kappa_{i+1})$. To perform encoding in the RSSE, the encoder performs the same steps as in the stochastic syndrome encoder in Section IV on $(\mathbf{H}_{i},\kappa_{i})$ for $i=1,2,\ldots$ (each with an independent common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$) until the first time where $\hat{\mathbf{V}}_{i}\neq\mathrm{e}$ (let it be iteration $L\in\mathbb{N}$), and selects $\hat{\mathbf{Z}}=\hat{\mathbf{V}}_{L}\boldsymbol{\Pi}_{L}^{-1}$, $\hat{\mathbf{Y}}=\mathbf{X}+\hat{\mathbf{Z}}$. The encoder then encodes $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$, and sends $(L,\mathbf{M})$. Practically, the encoder can encode $L$ using the Huffman code [46], Golomb code [47] or Elias code [48], and then append $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$ to form the transmitted bit sequence, which is prefix-free since the decoder can recover $L$ and then know the length of $\mathbf{M}$ which is $k_{L}$. The decoder can then decode $\mathbf{M}$ into $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$. If an optimal prefix-free code is used, the expected communication cost is at most $H(L)+\mathbb{E}[k_{L}]\log q+1$.
The precise algorithm for RSSE is given in Algorithm 1, where we assume there are two synchronized random number generators (RNGs) available to the encoder and the decoder which can be used to generate the common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$. In practice, they can initialize two pseudo-random number generators with the same seed.
@@
7:   $L\leftarrow i$
8:   $\hat{\mathbf{Z}}\leftarrow\hat{\mathbf{V}}_{L}\boldsymbol{\Pi}_{L}^{-1}$,  $\hat{\mathbf{Y}}\leftarrow\mathbf{X}+\hat{\mathbf{Z}}$
9:   Encode $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$
10:     (e.g., using (2) if $\mathbf{H}_{i}$ is in standard form)
11:   return $(L,\mathbf{M})$
12:  end if
13:end for
@@
    Output: output $\hat{\mathbf{Y}}\in\mathbb{F}_{q}^{n}$
@@
We now check that the RSSE simulates an AXN channel.
@@
The RSSE simulates an AXN channel, with noise distribution